OpenStudy (anonymous):

Simplify. Express the quotient as a rational exponent. x^4/5 / x^1/3 Simplify. Express the quotient as a rational exponent. x^4/5 / x^1/3 @Mathematics

OpenStudy (anonymous):

\[x ^{4/5} \div x ^{1/3}\] ^this is the question.

OpenStudy (agreene):

\[\huge x^{7/15}\]

OpenStudy (anonymous):

\[\Large \begin{array}{l} \frac{{{x^{4/5}}}}{{{x^{1/3}}}}\\ = {x^{(4/5) - (1/3)}}\\ = {x^{(12/15) - (5/15)}}\\ = {x^{7/15}} \end{array}\]